Nitrogen gas can be prepared by passing gaseous ammonia over solid copper (II) oxide at high temperatures. The other products of the reaction are solid copper and water vapour. If 18.10g of NH3 is reacted with 90.40g of CuO


1
Expert's answer
2021-05-26T07:07:56-0400

We start by balancing the chemical reaction:


"3\,CuO_{(s)}+2\,NH_{3(g)} \to N_{2(g)} +3\,Cu_{(s)}+3\,H_{2}O_{(l)}"


Then we use the amount of NH3 in grams and the proportions to find out the stoichiometry relation with the other reagent CuO:


"18.10g\,NH_3*\frac{1\,mol\,NH_3}{17.03\,g\,NH_3}*\frac{3\,mol\,CuO}{2\,mol\,NH_3}*\frac{79.55\,g\,CuO}{1\,mol\,CuO}"


"=126.82\,g\,CuO" (that will react entirely with 18.10g of NH3)


This means that CuO is the limiting reagent (because we only have 90.40 g of CuO at the start), thus we use that reagent mass to calculate the amount of N2 produced:


"90.40g\,CuO*\frac{1\,mol\,CuO}{79.55\,g\,CuO}*\frac{1\,mol\,N_2}{3\,mol\,CuO}*\frac{28.02\,g\,N_2}{1\,mol\,CuO}"


"=10.61\,g\,N_2\,produced"


In conclusion: CuO is the limiting reagent, NH3 is the reagent in excess and 10.61 g of N2 is produced while all the CuO is transformed into copper (Cu) and water by the reaction that was described with NH3.


Reference:

  • Chang, R. (2010). Chemistry. McGraw-Hill, New York.